# T2V-Turbo-v2: Enhancing Video Generation Model Post-Training through Data, Reward, and Conditional Guidance Design

### v2

@@ 4.3 Ablation Studies on the Design of Reward Models
Before investigating the impact of different selections of RMs, we can conclude that VCM falls short in aligning text and video from the results in Table 2. None of the VCM variants achieve a satisfactory Semantic Score, which motivates us to incorporate vision-language foundation models (Bommasani et al., 2021), such as CLIP and InternV2, to enhance text-video alignment in addition to feedback from HPSv2.1. In this section, we perform comprehensive ablation studies to assess the effectiveness of feedback from each model by training T2V-Turbo-v2 w/ MG with different combinations of HPSv2.1, CLIPScore, and InternV2 on VG + MV.
@@ 5 Related Work
Diffusion-based T2V Models. Conventional diffusion based T2V models often rely on large-scale image datasets for training (Ho et al., 2022b; Wang et al., 2023a; Chen et al., 2023; 2024a; Ho et al., 2022a) or adopt weights from pre-trained text-to-image (T2I) models (Zhang et al., 2023; Blattmann et al., 2023; Khachatryan et al., 2023). For example, LaVie (Wang et al., 2023c) begins training with WebVid-10M and LAION-5B before fine-tuning on a curated internal dataset of 23 million videos. Text-image datasets, such as LAION (Schuhmann et al., 2022), tend to be more than ten times larger than open-source video-text datasets like WebVid-10M (Bain et al., 2021), offering higher spatial resolution and greater diversity (Wang et al., 2023a). Recently, high-quality video datasets (Tan et al., 2024; Nan et al., 2024; Chen et al., 2024b; Wang et al., 2023d; Yang et al., 2024) with dense captions are collected and released to the public. In this paper, we aim to leverage the supervision signals from high-quality video to improve a pretrained T2V model during the post-training phase.
@@
Learning from Human/AI Feedback is an effective method for aligning generative models with human preferences (Leike et al., 2018; Ziegler et al., 2019; Ouyang et al., 2022; Stiennon et al., 2020; Rafailov et al., 2024; Li et al., 2024a). In the domain of image generation, various approaches have been introduced to align text-to-image models with human preferences, including reinforcement learning (RL)-based methods (Fan et al., 2024; Prabhudesai et al., 2023; Zhang et al., 2024) and backpropagation-based reward fine-tuning techniques (Clark et al., 2023; Xu et al., 2024; Prabhudesai et al., 2023). Recently, InstructVideo (Yuan et al., 2023) and Vader (Prabhudesai et al., 2024) extended reward fine-tuning to optimize text-to-video (T2V) models. Our method extends T2V-Turbo (Li et al., 2024b), integrating supervision from various domains, including high-quality datasets, diverse reward feedback, and conditional guidance.
Training-Free Conditional Guidance has been widely adopted in controlling T2I generations (Feng et al., 2023; Mo et al., 2024; Cao et al., 2023; Epstein et al., 2023; Ge et al., 2023; Patashnik et al., 2023; Xie et al., 2023) and achieves great success. MotionClone (Ling et al., 2024) tackles T2V generation by leveraging temporal attention from a reference video to guide the video generation process. In this paper, we leverage the same motion guidance by extracting motion prior from training videos and using them to formulate the energy function that augments the ODE solver. Our results show that this method enhances the motion quality in the generated videos.
@@ Appendix A Pseudo-codes of Our T2V-Turbo-v2’s Data Preprocessing and Training pipeline
for $(\boldsymbol{x},\boldsymbol{c})\in\mathcal{D}$ do
     Encode the video $\boldsymbol{x}$ into latent space: $\boldsymbol{z}=E(\boldsymbol{x})$
@@
     Sample $\boldsymbol{z}_{t_{n+k}}\sim\mathcal{N}\left(\alpha\left(t_{n+k}\right)\boldsymbol{z};\sigma^{2}\left(t_{n+k}\right)\mathbf{I}\right)$
     $\hat{\boldsymbol{z}}_{t_{n}}^{\Psi,\omega}\leftarrow\boldsymbol{z}_{t_{n+k}}+(1+\omega)\Psi(\boldsymbol{z}_{t_{n+k}},t_{n+k},t_{n},\boldsymbol{c};\psi)-\omega\Psi(\boldsymbol{z}_{t_{n}+k},t_{n+k},t_{n},\varnothing;\psi)$
@@
         Obtain the temporal attention $\mathcal{A}(\boldsymbol{z}_{t_{n+k}};\psi)$
         Perform DDIM inversion to obtain $\boldsymbol{z}^{\text{ref}}_{t_{n+k}}$
         Obtain the temporal attention $\mathcal{A}(\boldsymbol{z}^{\text{ref}}_{t_{n+k}};\psi)$
@@
         Calculate the motion guidance: $\mathcal{G}=\nabla_{\boldsymbol{z}_{t_{n+k}}}\left\|\mathcal{M}\circ\left(\mathcal{A}(\boldsymbol{z}_{t_{n+k}}^{\text{ref}})-\mathcal{A}(\boldsymbol{z}_{t_{n+k}})\right)\right\|_{2}^{2}$
@@
     Update $\mathcal{D}_{z}\leftarrow\mathcal{D}_{z}\cup\{\boldsymbol{z}_{t_{n+k}},\hat{\boldsymbol{z}}_{t_{n}}^{\Psi,\omega},\mathcal{G},\omega,\boldsymbol{c}\}$
@@
     Sample $(\boldsymbol{z}_{t_{n+k}},\hat{\boldsymbol{z}}_{t_{n}}^{\Psi,\omega},\mathcal{G},\omega,\boldsymbol{c})\sim\mathcal{D}_{z}$
     $\hat{\boldsymbol{z}}_{t_{n}}^{\Psi,\omega,\lambda}\leftarrow\hat{\boldsymbol{z}}_{t_{n}}^{\Psi,\omega}+\lambda\cdot\mathcal{G}$
     $\hat{\boldsymbol{x}}_{0}=\mathcal{D}\left(\boldsymbol{f}_{\theta}\left(\boldsymbol{z}_{t_{n+k}},\omega,\boldsymbol{c},t_{n+k}\right)\right)$
     $J_{\text{img}}(\theta)=\mathbb{E}_{\hat{\boldsymbol{x}}_{0},\boldsymbol{c}}\left[\sum^{M}_{m=1}\mathcal{R}_{\text{img}}\left(\hat{\boldsymbol{x}}^{m}_{0},\boldsymbol{c}\right)\right]$
     $J_{\text{vid}}(\theta)=\mathbb{E}_{\hat{\boldsymbol{x}}_{0},\boldsymbol{c}}\left[\mathcal{R}_{\text{v}}\left(\hat{\boldsymbol{x}}_{0},\boldsymbol{c}\right)\right]$
     $L_{\text{CD}}=d\left(\boldsymbol{f}_{\theta}\left(\boldsymbol{z}_{t_{n+k}},\omega,\boldsymbol{c},t_{n+k}\right),\texttt{stop\_grad}\left(\boldsymbol{f}_{\theta}\left(\hat{\boldsymbol{z}}_{t_{n}}^{\Psi,\omega},\omega,\boldsymbol{c},t_{n}\right)\right)\right)$
     $\mathcal{L}\left({\theta};\Psi\right)\leftarrow L_{\text{CD}}$ $-\beta_{\text{img}}J_{\text{img}}(\theta)-\beta_{\text{vid}}J_{\text{vid}}(\theta)$
     ${\theta}\leftarrow{\theta}-\eta\nabla_{{\theta}}\mathcal{L}\left({\theta},{\theta}^{-}\right)$
     ${\theta}^{-}\leftarrow\texttt{stop\_grad}\left(\mu{\theta}^{-}+(1-\mu){\theta}\right)$
@@ Appendix B Further Details about VBench
Total Score=4Quality Score+Semantic $\textbf{Total Score}=\frac{4\cdot\textbf{Quality Score}+\textbf{Semantic Score}}{5}$ (12)

### v3

@@ 4.3 Ablation Studies on the Design of Reward Models
From Table 2, we can conclude that VCM falls short in aligning text and video from the results. None of the VCM variants achieve a satisfactory Semantic Score, which motivates us to incorporate vision-language foundation models (Bommasani et al., 2021), such as CLIP and InternV2, to enhance text-video alignment in addition to feedback from HPSv2.1. In this section, we perform comprehensive ablation studies to assess the feedback from each model by training T2V-Turbo-v2 w/ MG with different combinations of HPSv2.1, CLIPScore, and InternV2 on VG + MV.
@@ 5 Related Work
Diffusion-based T2V Models. Conventional T2V models often rely on large-scale image datasets for training (Ho et al., 2022b; Wang et al., 2023a; Chen et al., 2023; 2024a; Ho et al., 2022a) or adopt weights from pre-trained text-to-image (T2I) models (Zhang et al., 2023; Blattmann et al., 2023; Khachatryan et al., 2023). For example, LaVie (Wang et al., 2023c) begins training with WebVid-10M and LAION-5B before fine-tuning on a curated internal dataset of 23 million videos. Text-image datasets, such as LAION (Schuhmann et al., 2022), tend to be more than ten times larger than open-source video-text datasets like WebVid-10M (Bain et al., 2021), offering higher spatial resolution and greater diversity (Wang et al., 2023a). Recently, high-quality video datasets (Tan et al., 2024; Nan et al., 2024; Chen et al., 2024b; Wang et al., 2023d; Yang et al., 2024) with dense captions are collected and released to the public. In this paper, we aim to leverage the supervision signals from high-quality video to improve a pretrained T2V model during the post-training phase.
@@
Learning from Human/AI Feedback is an effective method for aligning generative models with human preferences (Leike et al., 2018; Ziegler et al., 2019; Ouyang et al., 2022; Stiennon et al., 2020; Rafailov et al., 2024). In the domain of image generation, various approaches have been introduced to align T2I models with human preferences, including reinforcement learning (RL)-based methods (Fan et al., 2024; Prabhudesai et al., 2023; Zhang et al., 2024) and backpropagation-based reward fine-tuning techniques (Clark et al., 2023; Xu et al., 2024; Prabhudesai et al., 2023). Recently, InstructVideo (Yuan et al., 2023) and Vader (Prabhudesai et al., 2024) extended reward fine-tuning to optimize T2V models. Our method extends T2V-Turbo (Li et al., 2024b), integrating supervision from high-quality datasets, diverse reward feedback, and conditional guidance.
Training-Free Conditional Guidance has been widely adopted in controlling image generations (Feng et al., 2023; Mo et al., 2024; Cao et al., 2023; Epstein et al., 2023; Ge et al., 2023; Patashnik et al., 2023; Xie et al., 2023; Zhao et al., 2022) and achieves great success. MotionClone (Ling et al., 2024) tackles T2V generation by leveraging temporal attention from a reference video to guide the generation process. In this paper, we leverage the same motion guidance from training videos to augment the ODE solver, enhancing motion quality of the generated videos.
@@ 6 Conclusion and Limitations
Ethics Statement
In conducting this research, we have adhered to ethical guidelines to ensure integrity and transparency throughout the study. Our work does not involve human subjects, and therefore, no IRB approval was required. The datasets and models used in this study, such as OpenVid-1M, VigGen-1M, WebVid-10M, and VideoCrafter2, are publicly available and adhere to relevant data usage policies. We have taken care to use datasets that respect individual privacy, and no personally identifiable information has been included in our research.
Since we plan to release the code and models in the future, we are committed to formulating detailed guidelines to prevent the misuse of our models. These guidelines will emphasize ethical usage, discourage harmful applications, and promote responsible AI practices. Additionally, we will explore and incorporate safeguard mechanisms into our generation pipeline to mitigate risks associated with misuse, ensuring that the released models and code are aligned with ethical standards and safety considerations. This proactive approach is part of our ongoing responsibility to contribute to the safe and positive development of AI technologies.
Reproducibility Statement
Our experiments are conducted with all open-sourced codes and training data. Our implementation codes have been included in the supplementary material and will be released to the public in a GitHub repository without breaking the double-blind rules.
@@ Appendix A Pseudo-codes of Our T2V-Turbo-v2’s Data Preprocessing and Training pipeline
for $(\bm{x},\bm{c})\in\mathcal{D}$ do
  Encode the video $\bm{x}$ into latent space: $\bm{z}=E(\bm{x})$
@@
  Sample $\bm{z}_{t_{n+k}}\sim\mathcal{N}\left(\alpha\left(t_{n+k}\right)\bm{z};\sigma^{2}\left(t_{n+k}\right)\mathbf{I}\right)$
  $\hat{\bm{z}}_{t_{n}}^{\Psi,\omega}\leftarrow\bm{z}_{t_{n+k}}+(1+\omega)\Psi(\bm{z}_{t_{n+k}},t_{n+k},t_{n},\bm{c};\psi)-\omega\Psi(\bm{z}_{t_{n}+k},t_{n+k},t_{n},\varnothing;\psi)$
@@
   Obtain the temporal attention $\mathcal{A}(\bm{z}_{t_{n+k}};\psi)$
   Perform DDIM inversion to obtain $\bm{z}^{\text{ref}}_{t_{n+k}}$
   Obtain the temporal attention $\mathcal{A}(\bm{z}^{\text{ref}}_{t_{n+k}};\psi)$
@@
   Calculate the motion guidance: $\mathcal{G}=\nabla_{\bm{z}_{t_{n+k}}}\left\|\mathcal{M}\circ\left(\mathcal{A}(\bm{z}_{t_{n+k}}^{\text{ref}})-\mathcal{A}(\bm{z}_{t_{n+k}})\right)\right\|_{2}^{2}$
@@
  Update $\mathcal{D}_{z}\leftarrow\mathcal{D}_{z}\cup\{\bm{z}_{t_{n+k}},\hat{\bm{z}}_{t_{n}}^{\Psi,\omega},\mathcal{G},\omega,\bm{c}\}$
@@
  Sample $(\bm{z}_{t_{n+k}},\hat{\bm{z}}_{t_{n}}^{\Psi,\omega},\mathcal{G},\omega,\bm{c})\sim\mathcal{D}_{z}$
  $\hat{\bm{z}}_{t_{n}}^{\Psi,\omega,\lambda}\leftarrow\hat{\bm{z}}_{t_{n}}^{\Psi,\omega}+\lambda\cdot\mathcal{G}$
  $\hat{\bm{x}}_{0}=\mathcal{D}\left(\bm{f}_{\theta}\left(\bm{z}_{t_{n+k}},\omega,\bm{c},t_{n+k}\right)\right)$
  $J_{\text{img}}(\theta)=\mathbb{E}_{\hat{\bm{x}}_{0},\bm{c}}\left[\sum^{M}_{m=1}\mathcal{R}_{\text{img}}\left(\hat{\bm{x}}^{m}_{0},\bm{c}\right)\right]$
  $J_{\text{vid}}(\theta)=\mathbb{E}_{\hat{\bm{x}}_{0},\bm{c}}\left[\mathcal{R}_{\text{v}}\left(\hat{\bm{x}}_{0},\bm{c}\right)\right]$
  $L_{\text{CD}}=d\left(\bm{f}_{\theta}\left(\bm{z}_{t_{n+k}},\omega,\bm{c},t_{n+k}\right),\texttt{stop\_grad}\left(\bm{f}_{\theta}\left(\hat{\bm{z}}_{t_{n}}^{\Psi,\omega},\omega,\bm{c},t_{n}\right)\right)\right)$
  $\mathcal{L}\left({\theta};\Psi\right)\leftarrow L_{\text{CD}}$ $\beta_{\text{img}}J_{\text{img}}(\theta)-\beta_{\text{vid}}J_{\text{vid}}(\theta)$
  ${\theta}\leftarrow{\theta}-\eta\nabla_{{\theta}}\mathcal{L}\left({\theta};\Psi\right)$
@@ Appendix B Further Details about VBench
$\textbf{Total Score}=\frac{4\cdot\textbf{Quality Score}+\textbf{Semantic Score}}{5}$ (12)
@@ Appendix D Additional Qualitative Results
Appendix E Additional Ablation Studies on the design of training datasets
Table 12: Ablation studies on the design of training datasets. Based on the results in Table 2, we further conduct experiments on OV + VG and OV + VG + WV to corroborate the results.
VCM T2V-Turbo-v2 w/o MG
[…]
VG + WV
Quality Score 83.62 82.24 81.31 80.00 82.95 83.43 82.95 84.04 82.28 83.41 82.32 84.08 83.86 82.35
Semantic Score 61.93 58.06 55.51 46.53 60.65 57.52 54.98 68.73 72.22 73.04 75.74 78.33 69.25 76.80
Total Score 78.52 77.41 76.15 73.30 78.49 78.25 77.36 80.97 80.26 81.34 81.00 82.93 80.94 81.24
We further evaluate the performance of conduct experiments on the OV + VG and OV + VG + WV datasets to corroborate our conclusions in Table 2 in Sec. 4.2.
As shown in the table, VCM achieves a high Quality Score on the OV + VG dataset, similar to training on pure OV data, but adding the lower-quality WV data slightly decreases this score. Conversely, incorporating reward feedback in T2V-Turbo-v2 significantly improves the Semantic Score for the OV + VG + WV dataset, while the gains for OV + VG remain comparatively moderate. These findings align with our discussion in the main text: RMs with short context lengths operate optimally on datasets with shorter captions, highlighting the importance of aligning dataset characteristics with RM capabilities. Furthermore, the results justify our decision to exclude OV data when training the main T2V-Turbo-v2 models, as incorporating OV into VG + WV datasets negatively impacts model performance.
Appendix F Comparing Our T2V-Turbo-v2 with T2V-Turbo via Human Evaluation
Figure 9: Human evaluation on 16-step generation of T2V-Turbo-v2 w/o MG and T2V-Turbo-v2 w/ MG shows a clear preference for the latter. By incorporating motion information extracted from the training videos, T2V-Turbo-v2 w/ MG consistently produces videos that are favored for their superior motion quality and overall appeal.
We conduct a human evaluation to compare the 16-step video generation of T2V-Turbo-v2 w/o MG and T2V-Turbo-v2 w/ MG to verify the effectiveness of motion guidance. We carefully select 50 prompts that explicitly require motion generation from VBench. Each method generates 5 videos for each prompt with the same set of random seeds.
We hire annotators from Amazon Mechanical Turk to answer two questions: Q1) Which video demonstrates better motion quality? Q2) Which video do you prefer given the prompt?
We form the video comparison task as many batches of HITs. To ensure the annotation quality, we ensure the annotators are from English-speaking countries, including AU, CA, NZ, GB, and the US. Each task needs around 30 seconds to complete, and we pay each submitted HIT with 0.2 US dollars. Therefore, the hourly payment is about 24 US dollars. Note that the data annotation part of our project is classified as exempt by Human Subject Committee via IRB protocols.
The human evaluation results in Figure 9 show that videos generated by T2V-Turbo-v2 w/ MG are consistently preferred over those from T2V-Turbo w/o MG in terms of motion quality and overall appeal. These findings corroborate our automatic evaluation in Table 4, confirming that incorporating motion guidance significantly enhances model performance and improves the motion quality of generated videos.